OpenStudy (anonymous):

how to find derivative of function -x^3+2x^2+15x-1

OpenStudy (anonymous):

\[\frac{ d}{ dx }(f(x)) = -1\frac{ d }{ dx }(x^3)+2\frac{ d }{ dx }(2x^2)+15\frac{ d }{ dx }(x)-\frac{ d }{ dx }(1)\]

OpenStudy (anonymous):

use the power rule and the constant multiple rule to find the derivative

OpenStudy (anonymous):

thanks for your reply but i also need to find max and min valuse of the function

OpenStudy (anonymous):

oh, ok. I am not really sure how to find the min and max values then

OpenStudy (anonymous):

alright